Grade 11Physical Chemistry

Please Help..
How can i calculate the atomic mass of calcium if it is given 1.0395 gram of CaCl2 gives 2.6847 gram of AgCl?

Suhail Muhammed
11 Years agoGrade 11

1 Answer

Sunil Kumar FP
11 Years ago
IT can be calculated,let the atomic weight of Ca be x.molecular weight ofAgCl=143.3 and mol.wt of CaCl2=70+x
we have equlvalent of CaCL2=equivalent of AgCl
wt/eqwt Cacl2=wt/eq.wt of AgCl
1.0395/(70+x)/2=2.6847/143.3
x=40.9
Thus the atomic weight of Ca is 40(approx)